OpenStudy (anonymous):

the five consecutive intergers below add up to 390 x-3 x-2 x-1 x x+1 what is the value of x A 79 B 80 C 81 D 82 E 83

OpenStudy (anonymous):

uhhh idk

OpenStudy (anonymous):

part of the question

OpenStudy (arindameducationusc):

add the x's and add the numbers.... 5x+(-3-2-1+1)=390 =>5x+(-5)=390 =>5x=395 =>x=395/5 =79

OpenStudy (anonymous):

@arindameducationusc thx on how 2 solve it

OpenStudy (arindameducationusc):

No problem @jamesr @Astrophysics is also a cool guy..... He helps a lot.....

OpenStudy (anonymous):

thx 2 him 2